OpenStudy (anonymous):

2(x + 4) − (x + 1) = 10 3 1 2 0

OpenStudy (anonymous):

@dmezzullo help please

OpenStudy (anonymous):

dude can u help me please

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@amistre64 help me PLEASE

OpenStudy (dmezzullo):

2 times x+2 times 4-(x-1)=10 2x+8-(x-1)=10 -8 -8 2x-(x-1)=2 +1=+1 2x(x)=3 -2x -2x x=1

OpenStudy (anonymous):

thankyou but i needed u to explain it to me not just give me the answer

OpenStudy (dmezzullo):

Thats all the work wut else do u need?

OpenStudy (anonymous):

ooh i couldnt see that all i saw was x=1 thanks man

OpenStudy (dmezzullo):

Ya np

OpenStudy (dmezzullo):

SOrry it to so long i had to do something

OpenStudy (anonymous):

its ok